OpenStudy (anonymous):

Dirk can clean a playground in 5 hours. Geralyn can clean a playground in only 4 hours. Which of the following can be used to determine the amount of time it would take for Dirk and Geralyn to clean a playground together? a. 1 over 5 plus 1 over x equals 1 over 4 b. 1 over x plus 1 over 4 equals 1 over 5 c. 1 over 5 plus 1 over 4 equals 1 over x d. 1 over 4 plus 1 over 5 equals x over 9

jimthompson5910 (jim_thompson5910):

Hint: x = time it takes Dirk to do the job alone y = time it takes Geralyn to do the job alone z = time it takes for them to do the job working together \[\Large \frac{1}{x}+\frac{1}{y}=\frac{1}{z}\]

OpenStudy (anonymous):

@jim_thompson5910 thank you so so much!

jimthompson5910 (jim_thompson5910):

you're welcome